# Training a neural netwok for data reduction and better generalization

### v2

@@ 3 Monte Carlo simulation with regression task
We examine the performance of harderLASSO linear and artificial neural network by comparing them to various state-of-the-art models on the regression task that consists in predicting an output scalar $y$ from an input vector ${\bf x}$. The seven learners under consideration, all trained on the same datasets, are summarized in Table 1. All the LASSO-based methods, except LassoNet, use the square-root $\ell_{2}$-loss and select the penalty parameter $\lambda$ with the quantile universal rule (Giacobino et al., 2017b). Once $\lambda^{\rm QUT}$ value is calculated, the models are trained using the same $\lambda_{k}$ sequence as described in Section 2.5. LassoNet is a different neural network framework that achieves feature selection by adding a skip layer and integrating feature selection directly into the objective function; we select the tuning parameters of LassoNet using the built-in 10-fold cross-validation. Random Forest is based on the sklearn package with a configuration of 100 trees. The XGBoost learner is based on the xgboost Python package with default parameter settings. Given that both Random Forest and XGBoost primarily generate feature rankings rather than explicitly selecting features, we apply the Boruta algorithm to identify the most relevant features (Kursa et al., 2010). After determining the selected features using Boruta, we retrain the models on the selected subset: so random forest and XGBoost are two step methods here. If not stated otherwise, all neural network based models use one hidden layer of size $p_{2}=20$.
@@
harderLASSO linear Square-root $\ell_{2}$-loss and near hard thresholding
LASSO linear Square-root $\ell_{2}$-loss and $\ell_{1}$ penalty (Belloni et al., 2011)
@@
1. LASSO ANN Same as above (Ma et al., 2022)
2. LassoNet $\ell_{2}$-loss with mixed penalties (Lemhadri et al., 2021)
@@
3. Random Forest (RF) Average of CART (Breiman, 2001)
4. XGBoost Extreme Gradient Boosting (Chen and Guestrin, 2016)
@@ 3.3 Deeper ANN for a more complex nonlinearity
$W_{1}=\begin{pmatrix}1&-1&0&0\\ -1&1&0&0\\ 0&0&1&-1\\ 0&0&-1&1\\ \end{pmatrix},\quad W_{2}=\begin{pmatrix}10&10&-10&-10\\ -10&-10&10&10\\ \end{pmatrix}.$
@@
Although we have described the neural network using the ReLU activation function for the sake of generality, recall that in Definition 1 we specified that ReLU cannot be used within our neural networks. This difference is acceptable because employing a different activation function—such as a shifted version of ReLU or the ELU with $\alpha=1$—would result in the same network structure, albeit with different bias values. The key takeaway is that accurately representing the function requires two hidden layers.
@@
We investigate the performance of ANNs with 1, 2 and 3 hidden layers (with $p_{2}=20$, $(p_{2},p_{3})=(20,10)$ and $(p_{2},p_{3},p_{4})=(20,10,5)$) to fit the nonlinear association (16) using a Monte Carlo simulation. Here we choose $p_{1}=50$ inputs and generate a single $6500\times p_{1}$ input Gaussian matrix and selects its first $n$ rows to create data sets of size $n$ with increasing sample size ranging from 100 to 6500 by steps of 100. Then we simulate $m=50$ submatrices $X_{n,s}^{(k)}$ and generate $m$ training sets with (16), as in the previous simulation.
Figure 4 displays the Monte Carlo results. We observe that the 1 hidden layer ANN (blue) has a rapidly growing PESR (left), but its corresponding predictive performance (right) plateaus at a high level since it does not have enough flexibility to match the full nonlinearity of the association. The 2 hidden layer ANN (green) also has a rapid PESR increase, though less early, and its predictive performance does not plateau but keeps on improving. The 3 hidden layer ANN (yellow) has an even slower PESR increase and its predictive performance nearly matches that of the 2 hidden layer ANN when $n$ gets large.
@@ 4 Applications with classification task
USPS Images 9298 257 10 Kaggle
@@ 6 Reproducible research
The codes used to create the figures and tables are available at https://github.com/VcMaxouuu/AnnLasso.
7 Acknowledgement
We thank Dr. Thomas Kerdreux for his helpful comments. This work was partially supported by the Swiss National Science Foundation grant 200021E_213166 and the Academic Rising Star Project No. 2024-YJRC-LXY-005 of the National University of Defense and Technology, China.
Appendix A Proof of Threorem 1
@@
Appendix B Proof of Theorem 2
@@
$|{\cal L}_{n}({\cal Y},\mu_{{\bm{\theta}}^{\epsilon}}({\cal X}))-{\cal L}_{n}({\cal Y},\mu_{{\bm{\theta}}^{0}}({\cal X}))|$ $=$ $|\nabla_{{\bm{\theta}}_{1}}{\cal L}_{n}({\cal Y},\mu_{{\bm{\theta}}^{0}}(x))(\epsilon{\bm{\theta}}_{1})+\nabla_{{\bf c}_{1}}{\cal L}_{n}({\cal Y},\mu_{{\bm{\theta}}^{0}}({\cal X}))(\epsilon{\bf c}_{1})$
@@
${\cal L}_{n}({\cal Y},\mu_{{\bm{\theta}}^{\epsilon}}({\cal X}))+\lambda P_{\nu}(\epsilon{{\bm{\theta}}}^{(1)})$ $\geq{\cal L}_{n}({\cal Y},\mu_{{\bm{\theta}}^{0}}({\cal X}))-|\epsilon|\lVert g_{0}({\cal Y},{\cal X},{\bm{\theta}}^{0})\rVert_{\infty}+\lambda P_{\nu}(\epsilon{{\bm{\theta}}}^{(1)})+o(\epsilon^{2})$

### v3

@@ 3 Monte Carlo simulation with regression task
We examine the performance of harderLASSO linear and artificial neural network by comparing them to various state-of-the-art models on the regression task that consists in predicting an output scalar $y$ from an input vector ${\bf x}$. The seven learners under consideration, all trained on the same datasets, are summarized in Table 1. All the LASSO-based methods, except LassoNet, use the square-root $\ell_{2}$-loss and select the penalty parameter $\lambda$ with the quantile universal rule Giacobino et al. (2017b). Once $\lambda^{\rm QUT}$ value is calculated, the models are trained using the same $\lambda_{k}$ sequence as described in Section 2.5. LassoNet is a different neural network framework that achieves feature selection by adding a skip layer and integrating feature selection directly into the objective function; we select the tuning parameters of LassoNet using the built-in 10-fold cross-validation. Random Forest is based on the sklearn package with a configuration of 100 trees. The XGBoost learner is based on the xgboost Python package with default parameter settings. Given that both Random Forest and XGBoost primarily generate feature rankings rather than explicitly selecting features, we apply the Boruta algorithm to identify the most relevant features Kursa et al. (2010). After determining the selected features using Boruta, we retrain the models on the selected subset: so random forest and XGBoost are two step methods here. If not stated otherwise, all neural network based models use one hidden layer of size $p_{2}=20$.
@@
harderLASSO linear Square-root $\ell_{2}$-loss
and near hard thresholding
LASSO linear Square-root $\ell_{2}$-loss
and $\ell_{1}$ penalty Belloni et al. (2011)
@@
1. LASSO ANN Same as above Ma et al. (2022)
2. LassoNet $\ell_{2}$-loss with mixed penalties Lemhadri et al. (2021)
@@
3. Random Forest (RF) Average of CART Breiman (2001)
4. XGBoost Extreme Gradient Boosting Chen and Guestrin (2016)
@@ 3.3 Deeper ANN for a more complex nonlinearity
$W_{1}=\begin{pmatrix}1&-1&0&0\\ -1&1&0&0\\ 0&0&1&-1\\ 0&0&-1&1\\ \end{pmatrix},$
$W_{2}=\begin{pmatrix}10&10&-10&-10\\ -10&-10&10&10\\ \end{pmatrix}.$
@@
Although we have described the neural network using the ReLU activation function for the sake of generality, recall that in Definition 1.1 we specified that ReLU cannot be used within our neural networks. This difference is acceptable because employing a different activation function-such as a shifted version of ReLU or the ELU with $\alpha=1$-would result in the same network structure, albeit with different bias values. The key takeaway is that accurately representing the function requires two hidden layers.
@@
We investigate the performance of ANNs with 1, 2 and 3 hidden layers (with $p_{2}=20$, $(p_{2},p_{3})=(20,10)$ and $(p_{2},p_{3},p_{4})=(20,10,5)$) to fit the nonlinear association (16) using a Monte Carlo simulation. Here we choose $p_{1}=50$ inputs and generate a single $6500\times p_{1}$ input Gaussian matrix and selects its first $n$ rows to create data sets of size $n$ with increasing sample size ranging from 100 to 6500 by steps of 100. Then we simulate $m=100$ submatrices $X_{n,s}^{(k)}$ and generate $m$ training sets with (16), as in the previous simulation.
Figure 4 displays the Monte Carlo results. We observe that the 1 hidden layer ANN has a rapidly growing PESR (left), but its corresponding predictive performance (right) plateaus at a higher level since it does not have enough flexibility to match the full nonlinearity of the association. The 2 hidden layer ANN also has a rapid PESR increase, though less early, and its predictive performance does not plateau but keeps on improving eventually reaching the minimum. The 3 hidden layer ANN has an even slower PESR increase and its predictive performance nearly matches that of the 2 hidden layer ANN when $n$ gets large.
@@ 4 Applications with classification task
USPS Images 9298 256 10 Kaggle
@@ 6 Reproducible research
The codes used to create the figures and tables are available at https://github.com/VcMaxouuu/AnnHarderLasso.
Appendix A Proof of Threorem 2.4
@@
Appendix B Proof of Theorem 2.5
@@
$|{\cal L}_{n}({\cal Y},\mu_{{\bm{\theta}}^{\epsilon}}({\cal X}))-{\cal L}_{n}({\cal Y},\mu_{{\bm{\theta}}^{0}}({\cal X}))|$
$=|\nabla_{{\bm{\theta}}_{1}}{\cal L}_{n}({\cal Y},\mu_{{\bm{\theta}}^{0}}(x))(\epsilon{\bm{\theta}}_{1})+\nabla_{{\bf c}_{1}}{\cal L}_{n}({\cal Y},\mu_{{\bm{\theta}}^{0}}({\cal X}))(\epsilon{\bf c}_{1})$
@@
${\cal L}_{n}({\cal Y},\mu_{{\bm{\theta}}^{\epsilon}}({\cal X}))+\lambda P_{\nu}(\epsilon{{\bm{\theta}}}^{(1)})$
$\geq{\cal L}_{n}({\cal Y},\mu_{{\bm{\theta}}^{0}}({\cal X}))-|\epsilon|\lVert g_{0}({\cal Y},{\cal X},{\bm{\theta}}^{0})\rVert_{\infty}+\lambda P_{\nu}(\epsilon{{\bm{\theta}}}^{(1)})+o(\epsilon^{2})$